# Dynamics of Spontaneous Topic Changes in Next Token Prediction with Self-Attention

### v3

@@ Abstract
Human cognition is punctuated by abrupt, spontaneous shifts between topics—driven by emotional, contextual, or associative cues—a phenomenon known as spontaneous thought in neuroscience. In contrast, self-attention based models depend on structured patterns over their inputs to predict each next token, lacking spontaneity. Motivated by this distinction, we characterize spontaneous topic changes in self-attention architectures, revealing both their similarities and their divergences from spontaneous human thought. First, we establish theoretical results under a simplified, single-layer self-attention model with suitable conditions by defining the topic as a set of Token Priority Graphs (TPGs). Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a spontaneous topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, the longer context length or the more ambiguous input topic reduces the likelihood of spontaneous change. Second, we empirically validate that these dynamics persist in modern, state-of-the-art LLMs, underscoring a fundamental disparity between human cognition and AI behaviour in the context of spontaneous topic changes. To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human thought.
@@ 1 Introduction
Human cognition is punctuated by abrupt, apparently unstructured topic changes, the hallmark of human spontaneous thought, a phenomenon that has become a central topic in cognitive neuroscience (Bellana et al.,, 2022; Christoff and Fox,, 2018; Christoff et al.,, 2011; Kucyi et al.,, 2023; Mildner and Tamir,, 2019, 2024; Mills et al.,, 2020). For example, a spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling. These abrupt changes may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory. However, LLMs rely on contextual signals in the input to perform topic shifts, instead of spontaneous topic changes. They follow a structured, statistical approach, remaining on topic unless explicit cues signal a change. Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” (García Márquez,, 1967).
@@
Our work takes initial steps toward formalizing the dynamics of spontaneous topic changes in LLMs and analyzing how they relate to or diverge from human spontaneous thought. To this end, we ground our theoretical analysis in a single-layer self-attention model and empirically extend it to modern LLMs, laying groundwork for drawing parallels between AI models and human cognition.
@@
Recent advancements in the related field have substantially deepened our understanding of self-attention architectures. (Li et al., 2024b, ; Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ) have linked the self-attention to support vector machines (SVMs), offering optimization strategies for next-token prediction. A study by Li et al., (2023) highlights that in mixed-topics inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words. In parallel, prior studies have recognized the practical challenges of spontaneous topic changes in LLMs and proposed various approaches to address them (Hwang et al.,, 2024; Lim et al.,, 2010; Lin et al.,, 2023; Ni et al.,, 2022; Soni et al.,, 2022; Xie et al.,, 2021). Notably, spontaneous topic changes must be differentiated from hallucinations, generating incorrect or fabricated information without clear contextual basis (Ji et al.,, 2023; Maynez et al.,, 2020).
@@
Despite these advancements, our understanding of the dynamics of spontaneous topic changes in LLMs remains limited. Investigating the parallels and differences between spontaneous topic changes in self-attention models and human spontaneous thought could provide valuable insights for improving the efficiency of language models with human-like thought. Since modern LLMs rely on self-attention architectures, we begin by theoretically characterizing spontaneous topic shifts in a simplified setting. We then extend these findings through experiments on more complex, state-of-the-art models. To the best of our knowledge, no prior studies have investigated these dynamics so closely in relation to human thought.
@@
Figure 2 outlines our theoretical framework. To make the mathematical analysis tractable, we follow the same single-layer self-attention framework with log-loss objective function governed by Assumptions 14 from Li et al., 2024b . Inspired by token-priority graphs (TPGs) (Li et al., 2024b, ) and building on attribution graphs from Ameisen et al., (2025) for exposing an LLM’s internal computation, we define a topic as a set of TPGs. This graph-based formulation aligns naturally with recent advances in structured representations for LLMs (Sen et al.,, 2023; Wang et al.,, 2025). Furthermore, this mirrors neuroscience models of human spontaneous thought, in which concepts serve as nodes connected by associative edges (Mildner and Tamir,, 2019). Despite relying on these specific settings, our experiments extend our findings to modern LLMs, empirically confirming that relaxing these assumptions does not seem to undermine our core insights.
@@
Figure 1: Illustration of the difference between human cognition and LLMs. The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear spontaneous thought, but the GPT-2’s completion (bottom), demonstrates continuity. [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output.
@@
Figure 2: Overview of our theoretical framework. We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def. 2) and generate a dataset for each topic. The combination of dataset A and dataset B becomes the dataset for the mixed-topic model. We train self-attention models independently on each dataset. Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic). The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def. 3), ambiguous sequence (from Def. 4), changes topic A (change of topic from Def. 5). Further details for each category are in Figure 3.
Summary of Findings
@@
    Preservation of input topic priorities. Using a controlled sandbox, we demonstrate in Theorem 3 that attention models trained on mixed-topic datasets maintain the priorities of tokens associated with the original topic of an input sequence (Topic A in our analogy).
@@
    Changing topics triggered by token frequency. In Theorem 4, we show that only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A, the oracle’s responses may reflect a change of topic.
@@
    Impact of sequence length and topic ambiguity. Theorem 5 establishes that longer input sequences decrease the likelihood of changing topics. Furthermore, topic ambiguity acts as a stabilizing factor, not increasing the frequency of spontaneous topic changes.
@@
    Difference between LLMs and human cognition. In Section 6 we empirically extend Theorem 5 to modern deeper LLMs. Unlike human cognition, where extended discussions often encourage spontaneous thoughts and topic ambiguity promotes cognitive connections, our results highlight an opposite behavior in LLMs: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
Overview of the paper structure. We begin with the problem setup in Sec 2. Sec 3 introduces the definition of topic, and Sec 4 examines how self-attention models allocate the token priorities within the mixed-topics. In Sec 5, we establish the conditions under which a self-attention model induces spontaneous topic changes and show the dynamics of topic changes with longer input sequences or the presence of topic ambiguity. We then extend our analysis to frontier LLMs in Sec 6. Related work and discussion are provided in Secs 7 and 8, respectively. All proofs are provided in Appendix A.
@@ 2.1 Next Topic Prediction with Self-attention Model
In line with the approach presented by Tarzanagh et al., 2023b and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem. Given a vocabulary of size $K$ with an embedding matrix ${\bf E}=[{\bf e}_{1}\ {\bf e}_{2}\ \cdots\ {\bf e}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, we aim to predict the next token ID $y\in[K]$ based on an input sequence ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ with ${\bf x}_{i}\in{\bf E}$ for all $i\in[T]$. The training dataset, denoted as
@@
contains sequences of varying lengths $T_{i}$. In our notation ${\bf x}$ is the embedding vector corresponding to the token ID $x$, this is ${\bf x}={\bf e}_{x}$. For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in Tarzanagh et al., 2023b . The self-attention embedding output
@@
We keep the first two assumptions from Li et al., 2024b :
@@
Assumption 1 represents a variation of the weight-tying approach commonly used in language models (Press and Wolf,, 2017; Vaswani et al.,, 2017). Once training is complete, for a new input sequence ${\bf X}$, and a model characterized by $\bf W$, we predict the next token ID $\hat{y}_{\bf w}$ based on greedy decoding the probabilities from the softmax of the classification output
@@ 2.2 Token-priority Graph and Global Convergence of the Self-attention Model
Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary. $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$. They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},y)\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf e}_{y}$ to every token ${\bf x}\in{\bf X}$. TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens with equal priority. For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions. We use the same notation as Li et al., 2024b , given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
@@
For any two distinct nodes $i,j$ in the same TPG, they either satisfy $(i\Rightarrow j)$, $(j\Rightarrow i)$ or $(i\asymp j)$. Nodes in each $\mathcal{G}^{(k)}$ represent indices in $[K]$, and SCC structure supports the self-attention mechanism’s ability to assign priority within sequences based on the conditioning last token. Theorem 2 of Li et al., 2024b proved that under Assumptions 1 and 2, the self-attention model learned through Algo-GD converges to the solution of the following Support Vector Machine (SVM) defined by the TPGs of the underlying dataset DSET
@@
[…]
@@
This convergence implies that the model predicts the next token based on priorities obtained from the SCCs within the TPG relevant to the last token of the input sequence. Unlike the work in Li et al., 2024b , which considers multiple loss functions in subsequent results, we focus exclusively on a log-loss function in this work, leaving the exploration of other loss functions for future research. We add here another reasonable assumption that prevents the probabilities in Equation 1 to be equal for improbable numerical reasons, and we present our first lemma.
@@
Suppose conditions from Theorem 1 and Assumption 3 hold. Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\text{ iff }(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
This means that the tokens that maximize the probability in Equation 1 are all within the same SCC leading to the following definition:
@@ 3 Defining Topics
In order to answer our research questions regarding the dynamics of topic changes we need to define the concept of a topic. In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET. Therefore, inspired by Ameisen et al., (2025) that introduces attribution graphs to reveal the LLMs’ internal computational structure, we define a topic as a set of TPGs:
@@
Our graph-based formulation aligns with recent advances in structured representations of LLMs (Sen et al.,, 2023; Wang et al.,, 2025). Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in Li et al., 2024b . This leads to the following reasonable assumption:
@@
Detailed explanation is provided in Appendix B. This assumption enables the application of the results from Li et al., 2024b , with the concepts of topics and TPGs being used interchangeably.
@@
Remark. If we have two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, we can generate two different datasets DSETa and DSETb. The union of $\{\mathcal{G}^{(k)}_{a}\}_{k=1}^{K}$ for $\mathbb{T}_{a}$ and $\{\mathcal{G}^{(k)}_{b}\}_{k=1}^{K}$ for $\mathbb{T}_{b}$ forms $\{\mathcal{G}^{(k)}_{ab}\}_{k=1}^{K}$ for the mixed-topics, denoted as the TPGs for the combined DSETa and DSETb.

### v4

@@ Abstract
Human cognition is punctuated by abrupt, spontaneous shifts between topics—driven by emotional, contextual, or associative cues—a phenomenon known as spontaneous thought in neuroscience. In contrast, self-attention-based models rely on structured patterns over their inputs to predict each next token, lacking spontaneity. Motivated by this distinction, we characterize spontaneous topic changes in self-attention architectures and reveal divergences from spontaneous human thought. First, we establish theoretical results under a simplified, single-layer self-attention model with suitable conditions by defining a topic as a set of Token Priority Graphs (TPGs). Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a spontaneous topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, the longer context length or the more ambiguous input topic does not increase the likelihood of spontaneous change. Second, we empirically validate that the effect of input length or topic ambiguity persists in modern, state-of-the-art LLMs, underscoring a fundamental disparity between human cognition and AI behavior in the context of spontaneous topic changes. To the best of our knowledge, no prior work has explored these questions with a focus so closely aligned to human thought.
@@ 1 Introduction
Human cognition is punctuated by abrupt, apparently unstructured topic changes, the hallmark of spontaneous human thought, a phenomenon that has become a central topic in cognitive neuroscience [Bellana2022, spontaneousthought, Christoff2011, Kucyi2023, MILDNER2019763, judith2024why, outofblue]. For example, a spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling. These abrupt changes may be due to an emotional connection, such as recalling reading a book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory. However, LLMs shift topics in response to contextual cues in the input, rather than initiating spontaneous topic changes on their own. They follow a structured, statistical approach, remaining on topic unless explicit cues signal a change. Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” [marquez1967].
@@
Our work takes initial steps toward formalizing the dynamics of spontaneous topic changes in LLMs and analyzing how they relate to or diverge from spontaneous human thought. To this end, we ground our theoretical analysis in a single-layer self-attention model and empirically extend it to modern LLMs, laying the groundwork for drawing comparisons between AI models and human cognition.
@@
Recent advancements in the related field have substantially deepened our understanding of self-attention architectures. li2024mechanics, tarzanagh2023transformers, ataee2023max have linked the self-attention to support vector machines (SVMs), offering optimization strategies for next-token prediction. li2023transformers highlight that in mixed-topic inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words. In parallel, prior studies have recognized the practical challenges of spontaneous topic changes in LLMs and proposed approaches to address them [hwang-etal-2024-mp2d, Lim2010AST, multi-granularity2023, Ni2021RecentAI, soni-etal-2022-empirical, xie-etal-2021-tiage-benchmark]. Notably, spontaneous topic changes must be differentiated from hallucinations, generating incorrect or fabricated information without a clear contextual basis [ji2023survey, maynez-etal-2020-faithfulness].
@@
Despite these advancements, our understanding of the dynamics of spontaneous topic changes in LLMs remains limited. Investigating the relationship between spontaneous topic changes in self-attention models and spontaneous human thought can provide valuable insights into the cognitive discrepancies of current language models compared with humans. Since modern LLMs rely on self-attention architectures, we begin by theoretically characterizing spontaneous topic changes in a simplified setting. We then extend these findings through experiments on more complex, state-of-the-art models. To the best of our knowledge, no prior studies have investigated these dynamics so closely in relation to human thought.
@@
Figure 2 outlines our theoretical framework. To make the mathematical analysis tractable, we follow the same single-layer self-attention framework with log-loss objective function governed by Assumptions 14 from li2024mechanics. Inspired by token-priority graphs (TPGs) [li2024mechanics] and building on attribution graphs from ameisen2025circuit for exposing an LLM’s internal computation, we define a topic as a set of TPGs. This graph-based formulation aligns naturally with recent advances in structured representations for LLMs [sen-etal-2023-knowledge, wang2025model]. Furthermore, this mirrors neuroscience models of spontaneous human thought, in which concepts serve as nodes connected by associative edges [MILDNER2019763]. Despite relying on these specific settings, our experiments extend our findings to modern LLMs, empirically confirming that relaxing these assumptions does not seem to undermine our core insights.
@@
Figure 1: Illustration of the difference between human cognition and LLMs. The original fragment of “One hundred years of solitude" [marquez1967] (top) has a clear spontaneous thought, but the GPT-2’s completion (bottom), demonstrates continuity. [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output.
@@
Figure 2: Overview of our theoretical framework. We define a topic as a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def. 2) and generate a dataset for each topic. The combination of dataset A and dataset B becomes the dataset for the mixed-topic model. We train self-attention models independently on each dataset. Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic). The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def. 3); ambiguous sequence (from Def. 4); or changes topic A (change of topic from Def. 5). Further details for each category are shown in Figure 3.
1.1 Summary of findings
@@
    Preservation of input topic priorities. Using a controlled sandbox, we demonstrate in Theorem 2 that self-attention models trained on mixed-topic datasets maintain the priorities of tokens associated with the original topic of an input sequence (Topic A in our analogy).
@@
    Changing topics triggered by token frequency. In Theorem 3, we show that the oracle’s responses may reflect a change of topic only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A.
@@
    Impact of input length and topic ambiguity. Theorem 4 establishes that longer input sequences decrease the likelihood of changing topics. Furthermore, input topic ambiguity acts as a stabilizing factor, not increasing the frequency of spontaneous topic changes.
@@
    Difference between LLMs and human cognition. In Section 6 we empirically extend Theorem 4 to modern, deeper LLMs. Unlike human cognition, where extended discussions often encourage spontaneous thoughts and topic ambiguity promotes cognitive connections, our results highlight the opposite behavior in LLMs: neither longer prompts nor greater topic ambiguity appreciably increases the likelihood of a spontaneous topic change.
Overview of the paper structure. We begin with the problem setup in Sec 2. Sec 3 introduces the definition of topic, and Sec 4 examines how self-attention models allocate the token priorities within the mixed topics. In Sec 5, we establish the conditions under which a self-attention model induces spontaneous topic changes and show the dynamics of topic changes with longer input sequences or the presence of topic ambiguity. We then extend our analysis to frontier LLMs in Sec 6. Related work and discussion are provided in Secs 7 and 8, respectively. All proofs are provided in Appendix A.
@@ 2.1 Next topic prediction with self-attention model
In line with the approach presented by ataee2023max and li2024mechanics, we frame the next-token prediction task as a multi-class classification problem. Given a vocabulary of size $K$ with an embedding matrix ${\bf E}=[{\bf e}_{1}\ {\bf e}_{2}\ \cdots\ {\bf e}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, we aim to predict the next token ID $y\in[K]$ based on an input sequence ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ with ${\bf x}_{i}\in{\bf E}$ for all $i\in[T]$. The training dataset, denoted as
@@
contains sequences of varying lengths $T_{i}$. In our notation $\bf x$} is the embedding vector corresponding to the token ID $x$, this is ${\bf x}={\bf e}_{x}$. For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in ataee2023max. The self-attention embedding output
@@
We keep the first two assumptions from li2024mechanics:
@@
Assumption 1 represents a variation of the weight-tying approach commonly used in language models [press2016using, vaswani2017attention]. Once training is complete, for a new input sequence $\bf X$}, and a model characterized by $\bf W$, we predict the next token ID $\hat{y}_{\bf w}$ based on greedy decoding the probabilities from the softmax of the classification output
@@ 2.2 Token-priority graph and global convergence of the self-attention model
li2024mechanics defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary. $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$. They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},y)\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf e}_{y}$ to every token ${\bf x}\in{\bf X}$. TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens with equal priority. For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions. We use the same notation as li2024mechanics, given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
@@
For any two distinct nodes $i,j$ in the same TPG, they either satisfy $(i\Rightarrow j)$, $(j\Rightarrow i)$ or $(i\asymp j)$. Nodes in each $\mathcal{G}^{(k)}$ represent indices in $[K]$, and SCC structure supports the self-attention mechanism’s ability to assign priority within sequences based on the conditioning last token. Theorem 2 of li2024mechanics proved that under Assumptions 1 and 2, the self-attention model learned through Algo-GD converges to the solution of the following Support Vector Machine (SVM) defined by the TPGs of the underlying dataset DSET
@@
$\text{s.t.}\quad({\bf e}_{i}-{\bf e}_{j})^{\top}{\bf W}{\bf e}_{k}\begin{cases}=0,&\forall(i\asymp j)\in\mathcal{G}^{(k)}\\ \geq 1,&\forall(i\Rightarrow j)\in\mathcal{G}^{(k)}\end{cases}\forall k\in[K].$
@@
This convergence implies that the model predicts the next token based on priorities obtained from the SCCs within the TPG relevant to the last token of the input sequence. Unlike the work in li2024mechanics, which considers both hard retrieval and soft composition components and examines multiple loss functions in subsequent results, we focus exclusively on a log-loss function in this work, leaving the exploration of other loss functions for future research. Since the soft composition component is not required for our subsequent definitions and theoretical results, we concentrate solely on the hard retrieval component.
We add here another reasonable assumption that prevents the probabilities in Equation 1 from being equal due to improbable numerical reasons, and we present our first lemma.
@@
Suppose conditions from Theorem 1 and Assumption 3 hold. Consider an input sequence $\bf X$} from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\text{ iff }(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
This means that the tokens that maximize the probability for weights $\tilde{\bf W}$ in Equation 1 are all within the same SCC leading to the following definition:
@@ 3 Defining topics
In order to answer our research questions regarding the dynamics of topic changes we need to define the concept of a topic. In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET. Therefore, inspired by ameisen2025circuit that introduces attribution graphs to reveal the LLMs’ internal computational structure, we define a topic as a set of TPGs:
@@
Our graph-based formulation aligns with recent advances in structured representations of LLMs [sen-etal-2023-knowledge, wang2025model]. Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in li2024mechanics. This leads to the following reasonable assumption:
@@
Detailed explanation is provided in Appendix B. This assumption enables the application of the results from li2024mechanics, with the concepts of topics and TPGs being used interchangeably.
@@
Remark. Given two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, with corresponding datasets DSETa and DSETb, the union of $\{\mathcal{G}^{(k)}_{a}\}_{k=1}^{K}$ and $\{\mathcal{G}^{(k)}_{b}\}_{k=1}^{K}$ forms the TPGs for the mixed topics $\mathbb{T}_{ab}$, denoted by $\{\mathcal{G}^{(k)}_{ab}\}_{k=1}^{K}$.